# GHOST: Hallucination-Inducing Image Generation for Multimodal LLMs

### v1

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2024, 2023; Bai et al., 2025a; Team et al., 2025) have demonstrated impressive performance across a wide range of vision and language tasks, including image captioning, visual question answering, and multimodal reasoning. However, they suffer from a potentially critical failure mode: object hallucination, where the model incorrectly predicts the presence of an object that is not present in the image. Such errors raise concerns about the reliability of MLLMs in safety-sensitive applications and underscore the need for systematic methods to stress-test their visual robustness.
@@
Together, these components make GHOST scalable and effective: for instance, on Qwen2.5-VL (Bai et al., 2025a), GHOST generates 2,816 hallucination-inducing images out of 9,423 initial inputs, a 29% success rate–compared to 0.1% in DASH (Augustin et al., 2025). GHOST also generalizes to reasoning-based MLLMs like GLM-4.1V-Thinking (Team et al., 2025), where it shifts the model’s reasoning trajectory to justify the presence of a nonexistent object (Figure 7). The resulting images are natural and high quality, as demonstrated by the FID results in Table 6. Figure 2 shows qualitative examples, where each column compares an original image (top) with its GHOST-edited counterpart (bottom). The subtle semantic changes introduced by GHOST cause the model to hallucinate its presence. Human evaluation confirms that GHOST images remain object-free from a human perspective, with 89% of responses agreeing the target object is absent.
@@ 2 Related works
Systematic Failures in MLLMs. Recent works have introduced frameworks to automatically identify failures in MLLMs (Tong et al., 2024a, b; Hosseini et al., 2025). Tong et al. (2024b) highlight MLLMs’ struggles with basic visual understanding, attributing them to limitations in CLIP-based vision encoders. In contrast, Tong et al. (2024a) focuses on failures stemming from the language modality. Hosseini et al. (2025) show that MLLMs often over-rely on spurious visual cues, leading to hallucinations when such cues are present but the object itself is absent. A large body of work has specifically examined object hallucination (Li et al., 2023; Leng et al., 2024; Lovenia et al., 2023; Hu et al., 2023; Wang et al., 2023), typically proposing benchmarks and evaluation protocols to quantify the issue. However, these approaches rely on curated image sets, constraining analysis to a fixed set of visual scenarios. In contrast, our method actively generates failure images, uncovering novel and unanticipated visuals that are unlikely to appear in static datasets.
@@
DASH (Augustin et al., 2025) is the most comparable prior work, as it also incorporates MLLM feedback for image generation. However, it operates directly over diffusion latents and requires the MLLM, the diffusion model, the generated image, and an object detector to all remain in the optimization loop, making it both time and resource intensive and forcing the use of a distilled single-step diffusion model. Moreover, DASH ultimately retrieves similar images from the real dataset based on the generated ones. GHOST addresses these limitations with a decoupled and more efficient design. In Section 5 and Appendix B, we provide comparisons with DASH.
Robustness of Diffusion Representations. It is well-established that deep networks trained with empirical risk minimization (ERM) often rely on spurious correlations to make predictions (Sagawa et al., 2019; Kirichenko et al., 2022; Noohdani et al., 2024; Parast et al., 2025). Recent studies have also identified such correlations in CLIP (Radford et al., 2021b) and autoregressive vision-language models (Wang et al., 2024a; Varma et al., 2024; Kim et al., 2023; Ye et al., 2024; Zheng et al., 2024). In contrast, diffusion models have shown greater robustness in their learned representations, as they aim to approximate the underlying data distribution rather than optimizing for a specific downstream task (Li et al., 2024; Luo et al., 2024). Li et al. (2024) demonstrate that U-Net-based latent diffusion models (Rombach et al., 2022) are less prone to shortcut solutions compared to discriminative models. DEEM (Luo et al., 2024) further argues that vision encoders in MLLMs, often based solely on models like CLIP-ViT, can be improved using the generative feedback of diffusion models to align the semantic distributions of the image encoder. They showed this alignment enhances robustness to out-of-distribution inputs and reduces visual hallucinations. Motivated by these insights, we leverage diffusion models to probe vulnerabilities in MLLMs.
Adversarial Attacks on MLLMs. GHOST can be viewed as a form of adversarial generation, but we highlight key distinctions from prior work. Methods like AnyAttack (Zhang et al., 2025a) and AttackVLM (Zhao et al., 2023) operate in pixel space and preserve the overall image semantics, aiming for imperceptible perturbations. In contrast, GHOST inserts semantic-level misleading cues, plausible to humans yet triggering hallucinations in MLLMs. While adversarial examples are sometimes transferable across models, GHOST exhibits a different kind of transfer: cues crafted for one model generalize to others, pointing to shared failure modes and spurious biases. Our approach is closer in spirit to content-based adversarial attacks such as Chen et al. (2023), but differs in both setting (image classifiers vs. MLLMs) and objective (causing hallucination rather than misclassification).
@@ 4.1 Bridging Embedding Spaces
where $\mathbf{X}_{v}\in\mathcal{D}$. The training setup is illustrated in Figure 3b. For more details see Appendix C.
@@ 4.2 Optimization Objective
Absence of the target object semantics: To discourage $c$ from encoding the target object directly, we penalize its similarity to the CLIP embedding of the object’s textual descriptions. Rather than relying on a single prompt, we define a set of textual templates $\mathcal{T}_{clip}$ (e.g., “a photo of a tt”) to represent the target object (See Appendix E.2 for more details). We compute the expected cosine similarity between $c$ and the CLIP embeddings of these templates:
@@
To prevent overfitting to a single phrasing, we construct a diverse set of semantically equivalent query templates and randomly sample $\mathbf{X}_{q}$ from this set at each optimization step (The templates are provided in AppendixE.4).
@@ 4.3 Guided Diffusion
As shown in Figure 3c, instead of starting from pure noise, we begin the reverse process from a partially noised latent of the original image. This design encourages the final image to preserve the high-level structure of the input while still allowing for subtle semantic shifts that can induce hallucination. Concretely, we encode the original image $\mathbf{X}_{v}$ into its VAE latent representation, apply forward diffusion for $t$ steps to obtain $z_{t}$, and then condition the reverse denoising on $c$. The noise level $t$ controls the trade-off between preserving the original image structure and allowing space for misleading cues. To account for diffusion stochasticity, we allow up to $N=4$ generation attempts per image. More details are provided in Appendix D.
We use an open-vocabulary object detector, OWLv2 (Minderer et al., 2024), as a filtering step to verify that the generated image does not actually contain the target object. If OWLv2 detects the target object, we conservatively discard the sample. Only when the model hallucinates the object and OWLv2 confirms its absence do we consider the hallucination-inducing process successful. Figure 4 provides a qualitative example: as we optimize the CLIP embedding to increase the model’s confidence in answering “Yes,” the images gradually incorporate misleading visual features, in this case, vase-like structures. Note that this figure is illustrative only; we do not generate images at each step of optimization.
@@ 5 Experiments
In this section, we present both quantitative and qualitative results of GHOST on three open-source MLLMs. We also evaluate the transferability of GHOST-generated images across a diverse set of additional models.
@@ 5.1 Results
Models. We evaluate GHOST on three open-source MLLMs accessed via HuggingFace: Qwen2.5-VL-7B-Instruct (Bai et al., 2025a), LLaVA-v1.6-Mistral-7B (Liu et al., 2024), and GLM-4.1V-Thinking (Team et al., 2025).
Dataset. We select 10 target object classes from the COCO dataset (Lin et al., 2014). For each class, we begin by excluding all images containing the target object based on COCO annotations. Among the remaining negatives, we select the top 1,000 images with the highest CLIP similarity to the object name. This sorting identifies visually diverse yet semantically related contexts where hallucination is more likely to occur. In Appendix E.1, we also evaluate GHOST without CLIP sorting and show that our results remain consistent. To isolate GHOST-induced failures, we discard any image where the MLLM already predicts the target object before optimization.
Results. We present the GHOST results in Table 5. For each model, “Samples” refers to the number of images (out of the initial 10,000 COCO images) in which the model does not perceive the target object. We then report how many hallucination-inducing images GHOST successfully generates, followed by the success rate (i.e., number of hallucinations divided by the number of samples). We also include results from DASH (Augustin et al., 2025), applied to the COCO dataset. Despite operating on a much smaller image pool, GHOST discovers orders of magnitude more hallucination-inducing samples. For instance, GHOST identifies 2,816 successful cases for Qwen2.5-VL, while DASH-LLM and DASH-OPT combined find only 99. It is important to note that DASH was originally designed to search over ReLAION-5B (LAION, 2024), a massive web-scale dataset. Its goal is not to maximize hallucination count, but to uncover naturally occurring failure cases across large-scale image corpora. Our comparison on COCO is intended to provide a point of reference against an existing method that shares a similar objective: identifying systematic hallucination-inducing inputs. Further details about our DASH experiments are provided in Appendix B.
We also break down GHOST’s performance by object class in Figure 5. GHOST discovers a substantial number of hallucination cases for all classes. More details on the class-wise results of our method are provided in Appendix F.
Table 1: GHOST and DASH (Augustin et al., 2025) results on $CO$. “Samples” reflects the size of the input pool each method operates over.
@@
Image Quality. To evaluate the quality of the samples generated by GHOST, we conduct both qualitative and quantitative analyses. Figure 6 presents visual comparisons between GHOST and two baseline diffusion methods: Stable Diffusion v2.1 (Rombach et al., 2022), where we prompt the model using the COCO caption of the image, and Stable Diffusion unCLIP, where the diffusion is conditioned on the CLIP embedding of the image. GHOST preserves the semantics of the original image while introducing minimal visual artifacts. Table 6 reports Fréchet Inception Distance, FID scores (Heusel et al., 2017), evaluating each method along two axes: (i) realism, measured by FID against COCO validation images, and (ii) semantic fidelity, measured by FID against the initial images used in diffusion. While GHOST achieves comparable realism to baseline methods, it outperforms them in semantic preservation, confirming its advantage in maintaining visual identity. We observe higher FID scores for GLM hallucinations, likely due to the model’s reasoning nature. Since we optimize a “thinking” token rather than the final answer directly, the optimization path is less aligned with the hallucination objective, making it harder to induce consistent visual drift. Also more qualitative samples are provided in Appendix F.
@@
Human Evaluation. To evaluate whether GHOST images preserve the absence of the target object from a human perspective, we asked 40 peers to review the images, resulting in over 3,000 total votes. They were shown the final GHOST images and asked whether they could see the target object. On average, 89% of responses for images optimized w.r.t LLaVA-v1.6 and 86.3% for Qwen2.5-VL indicated that the object was not present. For reference, DASH (Augustin et al., 2025) reports that 5.2% of their images contained the object and 7.8% were ambiguous, implying that 87% of their samples were perceived as object-free. Additional details are provided in Appendix A.
@@ 5.2 Transferability
We evaluate how images generated by GHOST on one model transfer to others. In addition to the three source models, we include two larger open-source models: LLaMA3.2 11B (Meta, 2024) and Aya 32B (Dash et al., 2025), as well as two closed-source models: GPT-4o (OpenAI, 2024) and Gemini2.5-Flash (Comanici et al., 2025). As shown in Table 3, images optimized on Qwen2.5-VL achieve a 66.5% hallucination success rate on GPT-4o. These results suggest that different models share common failure patterns. More details are provided in Appendix F.
@@ 6.1 Effect of Hyperparameters
Effect of $\tau$. We study how the confidence threshold $\tau$ impacts GHOST’s success by varying $\tau\in\{0.5,0.6,0.7,0.8,0.9\}$ on 400 images from two objects (200 images each), while keeping the number of optimization steps fixed at 100 and using the same learning rate. A higher $\tau$ requires the optimized embedding to achieve higher $p(y^{*}\mid\mathbf{X}_{q},\Pi(c))$, making optimization harder. As expected, this reduces the number of images that meet the threshold within the allowed steps (see Appendix E.7). However, as shown in Figure 9, images generated from embeddings optimized with higher $\tau$ are more likely to induce hallucination. These results confirm that the confidence score $p(y^{*}\mid\mathbf{X}_{q},\Pi(c))$ serves as a meaningful proxy for the model’s belief in object presence, even though the optimization is decoupled from the diffusion process.
Effect of $\lambda_{\text{clip}}$. We evaluate the impact of $\lambda_{\text{clip}}$ using the same setup. As shown in Figure 9, increasing $\lambda_{\text{clip}}$ reduces the object detection rate per generated image. This confirms that stronger penalties on CLIP similarity discourage the diffusion model from generating images containing the object. Additional results are provided in Appendix E.6.
Effect of $\lambda_{reg}$. As shown in Appendix E.5, increasing $\lambda_{reg}\in\{1.0,1.5,2.0\}$ on GLM samples lowers FID scores, indicating that this term helps preserve visual realism.

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) (Liu et al., 2024; 2023a; Bai et al., 2025; Team et al., 2025) have demonstrated impressive performance across a wide range of vision and language tasks, including image captioning, visual question answering, and multimodal reasoning. However, they suffer from a potentially critical failure mode: object hallucination, where the model incorrectly predicts the presence of an object that is not present in the image. Such errors raise concerns about the reliability of MLLMs in safety-sensitive applications and underscore the need for systematic methods to stress-test their visual robustness.
@@
Together, these components make GHOST scalable and effective: for instance, on Qwen2.5-VL (Bai et al., 2025), GHOST generates 2,816 hallucination-inducing images out of 9,423 initial inputs, a 29% success rate–compared to 0.1% in DASH (Augustin et al., 2025). GHOST also generalizes to reasoning-based MLLMs like GLM-4.1V-Thinking (Team et al., 2025), where it shifts the model’s reasoning trajectory to justify the presence of a nonexistent object (Figure 7). The resulting images are natural and high quality, as demonstrated by the FID results in Table 6. Figure 2 shows qualitative examples, where each column compares an original image (top) with its GHOST-edited counterpart (bottom). The subtle semantic changes introduced by GHOST cause the model to hallucinate its presence. Human evaluation confirms that GHOST images remain object-free from a human perspective, with 89% of responses agreeing the target object is absent.
@@ 2 Related works
Systematic Failures in MLLMs. Recent works have introduced frameworks to automatically identify failures in MLLMs (Tong et al., 2024a; b; Hosseini et al., 2025). Tong et al. (2024b) highlight MLLMs’ struggles with basic visual understanding, attributing them to limitations in CLIP-based vision encoders. In contrast, Tong et al. (2024a) focuses on failures stemming from the language modality. Hosseini et al. (2025) show that MLLMs often over-rely on spurious visual cues, leading to hallucinations when such cues are present but the object itself is absent. A large body of work has specifically examined object hallucination (Li et al., 2023; Leng et al., 2024; Lovenia et al., 2023; Hu et al., 2023; Wang et al., 2023), typically proposing benchmarks and evaluation protocols to quantify the issue. However, these approaches rely on curated image sets, constraining analysis to a fixed set of visual scenarios. In contrast, our method actively generates failure images, uncovering novel and unanticipated visuals that are unlikely to appear in static datasets.
@@
DASH (Augustin et al., 2025) is the most comparable prior work, as it also incorporates MLLM feedback for image generation. However, it operates directly over diffusion latents and requires the MLLM, the diffusion model, the generated image, and an object detector to all remain in the optimization loop, making it both time and resource intensive and forcing the use of a distilled single-step diffusion model. Moreover, DASH ultimately retrieves similar images from the real dataset based on the generated ones. GHOST addresses these limitations with a decoupled and more efficient design. In Section 5 and Appendix C, we provide comparisons with DASH.
Robustness of Diffusion Representations. It is well-established that deep networks trained with empirical risk minimization (ERM) often rely on spurious correlations to make predictions (Sagawa et al., 2019; Kirichenko et al., 2022; Noohdani et al., 2024; Parast et al., 2025). Recent studies have also identified such correlations in CLIP (Radford et al., 2021b) and autoregressive vision-language models (Wang et al., 2024; Varma et al., 2024; Kim et al., 2023; Ye et al., 2024; Zheng et al., 2024). In contrast, diffusion models have shown greater robustness in their learned representations, as they aim to approximate the underlying data distribution rather than optimizing for a specific downstream task (Li et al., 2024; Luo et al., 2024). Li et al. (2024) demonstrate that U-Net-based latent diffusion models (Rombach et al., 2022) are less prone to shortcut solutions compared to discriminative models. DEEM (Luo et al., 2024) further argues that vision encoders in MLLMs, often based solely on models like CLIP-ViT, can be improved using the generative feedback of diffusion models to align the semantic distributions of the image encoder. They showed this alignment enhances robustness to out-of-distribution inputs and reduces visual hallucinations. Motivated by these insights, we leverage diffusion models to probe vulnerabilities in MLLMs.
Adversarial Attacks on MLLMs. GHOST can be viewed as a form of adversarial generation, but we highlight key distinctions from prior work. Methods like AnyAttack (Zhang et al., 2025) and AttackVLM (Zhao et al., 2023) operate in pixel space and preserve the overall image semantics, aiming for imperceptible perturbations. In contrast, GHOST inserts semantic-level misleading cues, plausible to humans yet triggering hallucinations in MLLMs. While adversarial examples are sometimes transferable across models, GHOST exhibits a different kind of transfer: cues crafted for one model generalize to others, pointing to shared failure modes and spurious biases. Our approach is closer in spirit to content-based adversarial attacks such as Chen et al. (2023), but differs in both setting (image classifiers vs. MLLMs) and objective (causing hallucination rather than misclassification).
@@ 4.1 Bridging Embedding Spaces
where $\mathbf{X}_{v}\in\mathcal{D}$. The training setup is illustrated in Figure 3b. For more details see Appendix D.
@@ 4.2 Optimization Objective
Absence of the target object semantics: To discourage $c$ from encoding the target object directly, we penalize its similarity to the CLIP embedding of the object’s textual descriptions. Rather than relying on a single prompt, we define a set of textual templates $\mathcal{T}_{clip}$ (e.g., “a photo of a tt”) to represent the target object (See Appendix F.2 for more details). We compute the expected cosine similarity between $c$ and the CLIP embeddings of these templates:
@@
To prevent overfitting to a single phrasing, we construct a diverse set of semantically equivalent query templates and randomly sample $\mathbf{X}_{q}$ from this set at each optimization step (The templates are provided in AppendixF.4).
@@ 4.3 Guided Diffusion
As shown in Figure 3c, instead of starting from pure noise, we begin the reverse process from a partially noised latent of the original image. This design encourages the final image to preserve the high-level structure of the input while still allowing for subtle semantic shifts that can induce hallucination. Concretely, we encode the original image $\mathbf{X}_{v}$ into its VAE latent representation, apply forward diffusion for $t$ steps to obtain $z_{t}$, and then condition the reverse denoising on $c$. The noise level $t$ controls the trade-off between preserving the original image structure and allowing space for misleading cues. To account for diffusion stochasticity, we allow up to $N=4$ generation attempts per image. More details are provided in Appendix E.
We use an open-vocabulary object detector, OWLv2 (Minderer et al., 2024), as a filtering step to verify that the generated image does not actually contain the target object. If OWLv2 detects the target object, we conservatively discard the sample. Only when the model hallucinates the object and OWLv2 confirms its absence do we consider the hallucination-inducing process successful. Figure 4 provides a qualitative example: as we optimize the CLIP embedding to increase the model’s confidence in answering “Yes,” the images gradually incorporate misleading visual features, in this case, vase-like structures. Note that this figure is illustrative only; we do not generate images at each step of optimization. See Appendix I for additional detector analyses.
@@ 5 Experiments
In this section, we present quantitative results of GHOST on three open-source MLLMs. We also evaluate the transferability of GHOST-generated images across a diverse set of additional models.
@@ 5.1 Results
Models. We evaluate GHOST on three open-source MLLMs accessed via HuggingFace: Qwen2.5-VL-7B-Instruct (Bai et al., 2025), LLaVA-v1.6-Mistral-7B (Liu et al., 2024), and GLM-4.1V-Thinking (Team et al., 2025).
Dataset. We select 10 target object classes from the COCO dataset (Lin et al., 2014). For each class, we begin by excluding all images containing the target object based on COCO annotations. Among the remaining negatives, we select the top 1,000 images with the highest CLIP similarity to the object name. This sorting identifies visually diverse yet semantically related contexts where hallucination is more likely to occur. In Appendix F.1, we also evaluate GHOST without CLIP sorting and show that our results remain consistent. To isolate GHOST-induced failures, we discard any image where the MLLM already predicts the target object before optimization. We further evaluate GHOST on ObjectNet (Barbu et al., 2019) to test its behavior on a different visual distribution; results are provided in Appendix G.
Results. We present the GHOST results in Table 5. For each model, “Samples” refers to the number of images (out of the initial 10,000 COCO images) in which the model does not perceive the target object. We then report how many hallucination-inducing images GHOST successfully generates, followed by the success rate (i.e., number of hallucinations divided by the number of samples). We also include results from DASH (Augustin et al., 2025), applied to the COCO dataset. Despite operating on a much smaller image pool, GHOST discovers orders of magnitude more hallucination-inducing samples. For instance, GHOST identifies 2,816 successful cases for Qwen2.5-VL, while DASH-LLM and DASH-OPT combined find only 99. It is important to note that DASH was originally designed to search over ReLAION-5B (LAION, 2024), a massive web-scale dataset. Its goal is not to maximize hallucination count, but to uncover naturally occurring failure cases across large-scale image corpora. Our comparison on COCO is intended to provide a point of reference against an existing method that shares a similar objective: identifying systematic hallucination-inducing inputs. Further details about our DASH experiments are provided in Appendix C. We also break down GHOST’s performance by object class in Figure 5. GHOST discovers a substantial number of hallucination cases for all classes. More details on the class-wise results of our method are provided in Appendix K.
Table 1: GHOST and DASH (Augustin et al., 2025) results on COCO. “Samples” reflects the size of the input pool each method operates over.
@@
Image Quality. To evaluate the quality of the samples generated by GHOST, we conduct both qualitative and quantitative analyses. Figure 6 presents visual comparisons between GHOST and two baseline diffusion methods: Stable Diffusion v2.1 (Rombach et al., 2022), where we prompt the model using the COCO caption of the image, and Stable Diffusion unCLIP, where the diffusion is conditioned on the CLIP embedding of the image. GHOST preserves the semantics of the original image while introducing minimal visual artifacts. Table 6 reports Fréchet Inception Distance, FID scores (Heusel et al., 2017), evaluating each method along two axes: (i) realism, measured by FID against COCO validation images, and (ii) semantic fidelity, measured by FID against the initial images used in diffusion. While GHOST achieves comparable realism to baseline methods, it outperforms them in semantic preservation, confirming its advantage in maintaining visual identity. We observe higher FID scores for GLM hallucinations, likely due to the model’s reasoning nature. Since we optimize a “thinking” token rather than the final answer directly, the optimization path is less aligned with the hallucination objective, making it harder to induce consistent visual drift. We further report SSIM (Wang et al., 2004) perceptual similarity results in Appendix F.9. Also more qualitative samples are provided in Appendix K.
@@
Human Evaluation. To evaluate whether GHOST images preserve the absence of the target object from a human perspective, we ask 40 peers to review the images, resulting in over 3,000 total votes. They see the final GHOST images and indicate whether they can see the target object. On average, 89% of responses for images optimized w.r.t LLaVA-v1.6 and 86.3% for Qwen2.5-VL indicate that the object is not present. For reference, DASH (Augustin et al., 2025) reports that 5.2% of their images contained the object and 7.8% were ambiguous, implying that 87% of their samples are perceived as object-free. We also evaluate the naturalness of GHOST images and find that they are rated comparably natural to the diffusion baseline. Additional details are provided in Appendix B.
Runtime Efficiency. GHOST requires $\sim$10 seconds per sample on a single A100 GPU, making it about $5\times$ faster than DASH on average, even when DASH uses a distilled diffusion model. Detailed runtimes are reported in Appendix F.8. We also report runtime measurements at higher image resolutions in Appendix F.10.
@@ 5.2 Transferability
We evaluate how images generated by GHOST on one model transfer to others. In addition to the three source models, we include two larger open-source models: LLaMA3.2 11B (Meta, 2024) and Aya 32B (Dash et al., 2025), as well as two closed-source models: GPT-4o (OpenAI, 2024) and Gemini2.5-Flash (Comanici et al., 2025). As shown in Table 3, images optimized on Qwen2.5-VL achieve a 66.5% hallucination success rate on GPT-4o. These results suggest that different models share common failure patterns. More details are provided in Appendix K.
@@ 6.1 Effect of Hyperparameters
Effect of $\tau$. We study how the confidence threshold $\tau$ impacts GHOST’s success by varying $\tau\in\{0.5,0.6,0.7,0.8,0.9\}$ on 400 images from two objects (200 images each), while keeping the number of optimization steps fixed at 100 and using the same learning rate. A higher $\tau$ requires the optimized embedding to achieve higher $p(y^{*}\mid\mathbf{X}_{q},\Pi(c))$, making optimization harder. As expected, this reduces the number of images that meet the threshold within the allowed steps (see Appendix F.7). However, as shown in Figure 9, images generated from embeddings optimized with higher $\tau$ are more likely to induce hallucination. These results confirm that the confidence score $p(y^{*}\mid\mathbf{X}_{q},\Pi(c))$ serves as a meaningful proxy for the model’s belief in object presence, even though the optimization is decoupled from the diffusion process.
Effect of $\lambda_{\text{clip}}$. We evaluate the impact of $\lambda_{\text{clip}}$ using the same setup. As shown in Figure 9, increasing $\lambda_{\text{clip}}$ reduces the object detection rate per generated image. This confirms that stronger penalties on CLIP similarity discourage the diffusion model from generating images containing the object. Additional results are provided in Appendix F.6.
Effect of $\lambda_{reg}$. As shown in Appendix F.5, increasing $\lambda_{reg}\in\{1.0,1.5,2.0\}$ on GLM samples lowers FID scores, indicating that this term helps preserve visual realism.